A ball is thrown at angle of 30° above the horizontal with a speed of 10 m/s. After 0.5 sec, find the horizontal component of its velocity.


1
Expert's answer
2021-06-21T05:52:03-0400

the horizontal component of its velocity "v _x = vcosθ = 10 \times cos 30 = 8.66 \frac{m}{sec}"

after 5 sec

Velocity "v = v_x + at = 8.66 + 1\times 0.5 = 9.16 \frac{m}{sec}"